OpenStudy (anonymous):

Jason wants to choose 9 players for his track team. There are 12 players to choose from. How many different teams can Jason make?

OpenStudy (anonymous):

12 choose 9 \[\left(\begin{matrix}12 \\ 9\end{matrix}\right)\]

OpenStudy (anonymous):

12 times 9?

OpenStudy (anonymous):

No you use this: http://en.wikipedia.org/wiki/Combination

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Have you taken computations and permutations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

108

OpenStudy (anonymous):

9 times 12

OpenStudy (raden):

use the combination formula : nCr = n!/(n-r)!r! with n = 12 and r = 9